OpenStudy (anonymous):

What number would you have to add to both sides to complete the square? x(2) – 14x = 7

OpenStudy (anonymous):

what is half of 14?

OpenStudy (anonymous):

that is not the answer, that is step one

OpenStudy (anonymous):

7

OpenStudy (anonymous):

find half of 14 and then square it. That's what you add to both sides

OpenStudy (anonymous):

so it would be 196

OpenStudy (mathstudent55):

To complete the square, you must have an equation of the form x^2 + bx + c = 0, which you have. Just do what was written above. Divide the coefficient of the x-term by 2. Then square that.

OpenStudy (anonymous):

no \(7^2=49\)

OpenStudy (mathstudent55):

You need the square of 7, not of 14.

OpenStudy (anonymous):

ok so the answer is 49

OpenStudy (mathstudent55):

If the equation has the form ax^2 + bx + c = 0, where a is not 1, then you need to first divide both sides by a. For example, if you had 2x^2 + 24x + 4 = 0 Then you would first divide the entire equation by 2, the coefficient of the x^2 term. Then you can continue with dividing the x-term coefficient by 2 and squaring it.

OpenStudy (mathstudent55):

Yes, in this case, the answer is 49.

OpenStudy (anonymous):

thankyou mathstudent55